# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v2

@@
Abstract
Foundation Models (FMs) have improved time series forecasting in various sectors, such as finance, but their vulnerability to input disturbances can hinder their adoption by stakeholders, such as, investors and analysts. To address this, we propose a causally grounded rating framework to study the robustness of Foundational Models for Time Series (FMTS) with respect to input perturbations. We evaluate our approach on the stock price prediction problem, a well studied problem with easily accessible public data, evaluating six state-of-the-art (some multi-modal) FMTS across six prominent stocks spanning three industries. The ratings proposed by our framework effectively assess the robustness of FMTS and also offer actionable insights for model selection and deployment. Within the scope of our study, we find that (1) multi-modal FMTS exhibit better robustness and accuracy compared to their uni-modal versions and, (2) FMTS pre-trained on time series forecasting task exhibit better robustness and forecasting accuracy compared to general-purpose FMTS pre-trained across diverse settings. Further, to validate our framework’s usability, we conduct a user study showcasing FMTS prediction errors along with our computed ratings. The study confirmed that our ratings reduced the difficulty for users in comparing the robustness of different systems.
1 Introduction
@@
Time series (TS) forecasting uses historical data indexed by time to predict future values. This task finds wide applicability in industry in domains like finance, healthcare, manufacturing, and weather. Although well-studied, the TS forecasting has seen recent advancements with new AI-based approaches including gradient boosting, deep learning, transformers and Foundation Models (FMs) trained on uni-modal numerical data as well as multi-modal data vying for state-of-the-art performance (Elsayed et al. 2021; Jin et al. 2023).
@@
However, having good performance is no guarantee that users will trust a method or model and use it. In particular, users care about the model’s robustness to noisy data and perturbations, as erroneous predictions can have far-reaching impact on stakeholders. The perturbations may have been caused unintentionally by an actor or intentionally by an adversary, but regardless, the users expect robust and consistent performance. To manage user trust, a promising idea is of third-party assessment of models and ratings (automated certifications), which can help users make informed decisions, even without access to the method’s code or model’s training data using both statistical (Srivastava and Rossi 2018, 2020; Srivastava et al. 2024) and causality-based methods (Lakkaraju et al. 2023; Lakkaraju, Srivastava, and Valtorta 2024).
In this context, our contributions are that we: (a) introduce a novel workflow for assessing and rating FMTS for robustness through causal analysis. (b) introduce three perturbations for both numerical and line plots (image) data inspired by real-world applications in unintended scenarios. (c) introduce two novel metrics to measure the causal impact of other attributes on the FMTS (along with existing metrics from literature). (d) create ratings to compare the models in terms of forecasting accuracy and robustness. (e) conduct a user study to assess the ease of interpreting FMTS behavior through our ratings and to assess its alignment with users’ perceptions. With our causally grounded rating framework, we evaluate leading FMTS models—two general-purpose-data trained (Gemini-V and Phi-3 in both uni-modal and multi-modal forms) and two time series-data trained (MOMENT and Chronos) - across diverse architectures (encoder-only, decoder-only, and encoder-decoder) and parameter sizes (46M to 32B), along with three baseline models (ARIMA, random, and biased) (total of 9 models). We use one year of stock price data from six leading companies across three different industries as test data. We find that for both Gemini and Phi-3, their multi-modal versions, in general, exhibit better robustness and forecasting accuracy compared to their uni-modal versions (Figure 5). We also find that time series-specific FMTS exhibit better robustness and forecasting accuracy compared to general-purpose FMTS (Figure 6). Furthermore, the user study confirms that our ratings reduced the difficulty for users in comparing the robustness of different systems.
2 Related Work
@@
Foundation Models Supporting Time Series The use of FMs for time series forecasting has advanced significantly. (Lu et al. 2022) showed that transformers pre-trained on text data can solve sequence modeling tasks in other modalities, enabling their application to time series analysis. Recent studies have reprogrammed LLMs for time series tasks through parameter-efficient fine-tuning and tokenization strategies (Zhou et al. 2023; Gruver et al. 2023; Jin et al. 2023; Cao et al. 2023; Ekambaram et al. 2024). (Zhou et al. 2023) and (Jin et al. 2023) further illustrate the versatility and robustness of fine-tuned language pre-trained transformers for diverse time series tasks. Several models have advanced time series forecasting. (Ansari et al. 2024) and (Woo et al. 2024) have improved forecasting accuracy and model generalization, while (Rasul et al. 2023) and (Das et al. 2023) have explored new tokenization strategies and fine-tuning methods. (Garza and Mergenthaler-Canseco 2023) and (Ekambaram et al. 2024) developed lightweight models for real-time applications, and (Talukder, Yue, and Gkioxari 2024) integrated multiple temporal patterns to improve precision. FMs trained from scratch, like (Gruver et al. 2023), achieved SOTA in zero-shot forecasting, with (Cao et al. 2023) and (Goswami et al. 2024) further improving model performance. Please see Section 5.1 for the FMTS we selected due to their SOTA performance in their respective categories.
@@
Perturbations in Time Series Data TS data is commonly stored in spreadsheets and databases, which are prone to changes due to acts of omission (e.g., negligence, data-entry errors) or commission (e.g., adversarial attacks, sabotage). Omission errors are most common (Panko and Halverson 1996). Tools like Microsoft Excel and Google Sheets are widely used for data collection and analysis, allowing end-user programming (Birch, Lyford-Smith, and Guo 2018). However, over 90% of spreadsheets contain errors due to issues like incorrect formulae, leading to multi-billion dollar losses (Pak-Lok POON and TANG 2024). Adversarial attacks are also increasing in data stores and AI models for tasks like forecasting. (Karim et al. 2019) adapted these concepts to time series, exploring both black-box and white-box attacks. (Oregi et al. 2018) revealed the vulnerability of distance-based classifiers. (Rathore et al. 2020) examined various adversarial attacks on time series classifiers. TSFool (Li et al. 2022) introduced a multi-objective black-box attack to craft imperceptible adversarial time series to fool RNN classifiers.
@@
Rating AI Systems Several works have assessed and rated AI systems for trustworthiness from a third-party perspective without access to training data. (Srivastava and Rossi 2020) proposed a method to rate AI systems for bias, specifically targeting gender bias in machine translators (Srivastava and Rossi 2018), and used visualizations to communicate these ratings (Bernagozzi et al. 2021b). They conducted user studies on trust perception through visualizations (Bernagozzi et al. 2021a), but these lacked causal interpretation. (Lakkaraju, Srivastava, and Valtorta 2024) introduced a causal analysis approach to rate bias in sentiment analysis systems, extending it to assess their impact when used with translators (Lakkaraju et al. 2023). We extend their method to rate MM-TSFM for robustness against perturbations. Causal analysis offers advantages over statistical analysis by determining accountability, aligning with humanistic values, and quantifying the direct influence of various attributes on forecasting accuracy.
3 Problem
3.1 Preliminaries
@@
Time Series Forecasting Let the time series be represented by {xt-n+1, xt-n+2, …., xt, xt+1, …, xt+d}, where each xt-n+i represents a value in time series, where $n$ is called the sliding window size and $d$ is the number of future values the model predicts. Let Xt === {xt-n+1, xt-n+2, …., xt}, and $\hat{Y_{t}}$ $=$ $\hat{x}$t+1, $\hat{x}$t+2, …., $\hat{x}$t+d}, where $\hat{Y_{t}}=f(X_{t}$; $\theta$) for uni-modal FMTS, and in the case of multi-modal FMTS, $X_{t}$ includes both the numerical time series values and the corresponding time series line plots (images). The function $f$ represents pre-trained FMTS with parameter $\theta$ used in a zero-shot manner without task-specific fine-tuning that predicts the values for the next ‘d’ timesteps based on the values at previous $n$ timesteps. Let $Y_{t}$ denote the true values for the next ‘d’ timesteps. Let $S$ be the set of FMTS we want to rate. Let Rt be the residual for the sliding window [t + 1, t + d] and is computed by ($\hat{Y_{t}}$ $Y_{t}$) at each timestep. Our rating method aims to highlight the worst-case scenario for the model. Therefore, we consider the maximum residual, denoted as $R^{max}_{t}$.
Figure 1: Causal model $\mathcal{M}$ for FMTS. The validity of link ‘1’ depends on the data distribution ($P|Z$), while the validity of the links ‘2’ and ‘3’ are tested in our experiments.
@@
Figure 2: Variants of the causal diagram in Figure 1 used to answer different research questions (RQs).
@@
Causal Model The causal model $\mathcal{M}$, is shown in Figure 1. Arrowheads indicate the causal direction from cause to effect. If Sensitive Attribute ($Z$) is a common cause for both Perturbation ($P$) and Residual ($R^{max}_{t}$), it introduces a spurious correlation between $P$ and $R^{max}_{t}$, known as the confounding effect, making $Z$ the confounder. The path from Perturbation to Residual through the confounder is called a backdoor path and is undesirable. Various backdoor adjustment techniques can remove the confounding effect (Xu and Gretton 2022; Fang et al. 2024; Liu et al. 2021). The deconfounded distribution, after adjustment, is represented as $(R^{max}_{t}|do(P))$. The ‘do(.)’ operator in causal inference denotes an intervention to measure the causal effect of $P$ on the $R^{max}_{t}$. Solid red arrows with ‘?’ in Figure 1 denote the causal links tested in our experiments, while the dotted red arrow represents a potential causal link, depending on the distribution $(P|Z)$ across different values of $Z$.
3.2 Problem Formulation
@@
We aim to answer the following research questions (RQs) (with causal diagrams in Fig 2) through our causal analysis when different perturbations denoted by $P$ = {0, 1, 2, 3} (or simply $P0,P1,P2,P3$) are applied to the input given to the set of FMTS $S$:
@@
RQ1: Does $Z$ affect $R^{max}_{t}$, even though $Z$ has no effect on $P$? That is, if perturbations are independent of the sensitive attribute, can the attribute still affect the system outcome, leading to statistical bias (i.e., lack of fairness)? Causal analysis is unnecessary here due to no confounding effect.
@@
RQ2: Does $Z$ affect the relationship between $P$ and $R^{max}_{t}$ when $Z$ has an effect on $P$? That is, if the applied perturbations depend on the value of the sensitive attribute, would the sensitive attribute add a spurious (false) correlation between the perturbation and the outcome of a system leading to confounding bias?
@@
RQ3: Does $P$ affect $R^{max}_{t}$ when $Z$ may have an effect on $R^{max}_{t}$? That is, what is the impact of the perturbation on the outcome of a system when the sensitive attribute may still have an effect on the outcome of a system?
@@
RQ4: Does $P$ affect the accuracy of $S$? That is, do the perturbations affect the performance of the systems’ accuracy? Causal analysis is not required to answer this question as we only need to compute appropriate accuracy metrics to assess how robust a system is against different perturbations.
4 Solution Approach

### v3

# Creating a Causally Grounded Rating Method for Assessing the Robustness of AI Models for Time-Series Forecasting

@@
Abstract.
AI models, including both time-series-specific and general-purpose Foundation Models (FMs), have demonstrated strong potential in time-series forecasting across sectors like finance. However, these models are highly sensitive to input perturbations, which can lead to prediction errors and undermine trust among stakeholders, including investors and analysts. To address this challenge, we propose a causally grounded rating framework to systematically evaluate model robustness by analyzing statistical and confounding biases under various noisy and erroneous input scenarios. Our framework is applied to a large-scale experimental setup involving stock price data from multiple industries and evaluates both uni-modal and multi-modal models, including Vision Transformer-based (ViT) models and FMs. We introduce six types of input perturbations and twelve data distributions to assess model performance. Results indicate that multi-modal and time-series-specific FMs demonstrate greater robustness and accuracy compared to general-purpose models. Further, to validate our framework’s usability, we conduct a user study showcasing time-series models’ prediction errors along with our computed ratings. The study confirms that our ratings reduce the difficulty for users in comparing the robustness of different models. Our findings can help stakeholders understand model behaviors in terms of robustness and accuracy for better decision-making even without access to the model weights and training data, i.e., black-box settings.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
@@
Time series (TS) forecasting uses historical data indexed by time to predict future values. This task finds wide applicability in industry in domains like finance, healthcare, manufacturing, and weather. Although well-studied, the TS forecasting has seen recent advancements with new AI-based approaches including gradient boosting, deep learning, transformers and Foundation Models (FMs) trained on uni-modal numerical data as well as multi-modal data vying for state-of-the-art performance (Elsayed et al., 2021; Jin et al., 2023).
@@
However, having good performance is no guarantee that users will trust a method or model and use it. In particular, users care about the model’s robustness to noisy data and perturbations, as erroneous predictions can have far-reaching impact on stakeholders. The perturbations may have been caused unintentionally by an actor or intentionally by an adversary, but regardless, the users expect robust and consistent performance. To manage user trust, a promising idea is of third-party assessment of models and ratings (automated certifications), which can help users make informed decisions, even without access to the method’s code or model’s training data using both statistical (Srivastava and Rossi, 2018, 2020; Srivastava et al., 2024) and causality-based methods (Lakkaraju et al., 2023, 2024).
In this context, our contributions are that we:
  1. (1)
    We propose a rating workflow for rating time-series forecasting models (TSFM) for robustness, extending previous rating methods (Srivastava and Rossi, 2018, 2020; Srivastava et al., 2024; Lakkaraju et al., 2023, 2024). This workflow supports a new use case of model selection for time-series forecasting based on robustness and forecasting accuracy.
  2. (2)
    We define six input perturbations for evaluation: Input-specific Perturbations (IP) (2), Semantic Perturbations (SP) (2), Syntactic Perturbation (SyP) (1), and Composite Perturbation (CP) (1), which includes a joint evaluation with an image-based sentiment analysis system. We evaluate these models using one year of stock price data from six leading companies across three industries.
  3. (3)
    Our rating method introduces two novel causality-based metrics alongside established ones, generating ratings to compare three baseline models along with 11 time-series forecasting models (which include 9 foundation models and 2 ViT-num-spec models) on forecasting accuracy and robustness.
  4. (4)
    We conduct a user study to evaluate the usability of ratings to interpret model behavior. The study confirms that our ratings reduce the difficulty for users in comparing the robustness of different models.
  5. (5)
    In addition to the core research questions outlined in Section 3, through our experiments, we also answer additional research questions (ARQs) in Section 5.3:
    • ARQ1: How do different model types compare in performance? (Foundation Vs. Non-foundation models)
    • ARQ2: Does multi-modality improve the performance of TSFM? (uni-modal vs. multi-modal)
    • ARQ3: How does the architecture of FMs influence performance? (time series-specific vs. general-purpose and encoder-only vs. decoder-only vs. encoder-decoder)
2. Related Work
We now contextualize our work with related literature so that our contributions are highlighted. We cover Transformer-based TSFM models, perturbations in finance domain, robustness testing of TSFM, causal analysis of TSFM, and rating AI models.
2.1. Time-series Forecasting with Transformer-based Models
Transformer architectures have gained significant traction in time-series forecasting due to their effectiveness in capturing long-term temporal patterns and handling variable input lengths. (Lu et al., 2022) shows that transformers pre-trained on text data can solve sequence modeling tasks in other modalities paving the way for leveraging language pre-trained transformers for time series analysis. (Zhou et al., 2023) and (Jin et al., 2023) further illustrate the versatility and robustness of fine-tuned language pre-trained transformers for diverse time series tasks. (Cheng et al., 2022) introduces a multi-modal graph neural network for learning from multi-modal inputs. These works typically use data from various sources. In contrast, (Zeng et al., 2023) introduces a vision transformer using time-frequency spectrograms to transform numerical data into a multi-modal form, showing benefits in both time and frequency domains. We extend this work by evaluating two variations of the ViT-num-spec model from (Zeng et al., 2023). These works typically use data from multiple sources to support multi-modality or operate solely on single-modality data such as raw numerical sequences. In contrast, (Zeng et al., 2023) transforms simple numerical time-series data into a multi-modal representation using time-frequency spectrograms, enabling a vision transformer to jointly learn from both time intensities and frequency spectrograms. We extend this work by evaluating two variants of the ViT-num-spec model proposed in (Zeng et al., 2023).
@@ 2.1. Time-series Forecasting with Transformer-based Models
Building on the strengths of transformers, recent developments have introduced FMs that extend these capabilities to multi-task settings, enabling zero-shot forecasting and broader generalization across other domains including time-series. Recent studies have reprogrammed LLMs for time series tasks through parameter-efficient fine-tuning and tokenization strategies (Zhou et al., 2023; Gruver et al., 2023; Jin et al., 2023; Cao et al., 2023; Ekambaram et al., 2024). (Ansari et al., 2024) and (Woo et al., 2024) have improved forecasting accuracy and model generalization, while (Rasul et al., 2023) and (Das et al., 2023) have explored new tokenization strategies and fine-tuning methods. (Yu et al., 2023) leverages historical stock prices, company metadata, and economic news for explainable time series forecasting with LLMs. (Garza and Mergenthaler-Canseco, 2023) and (Ekambaram et al., 2024) developed lightweight models for real-time applications, and (Talukder et al., 2024) integrated multiple temporal patterns to improve precision. FMs trained from scratch, like (Gruver et al., 2023), achieved SOTA in zero-shot forecasting, with (Cao et al., 2023) and (Goswami et al., 2024) further improving model performance. In our experiments, we select Gemini-V and Phi-3 as the General Purpose FMs and Chronos and MOMENT as Time-series-specific FMs due to their SOTA performance in their respective categories.
2.2. Perturbations in Finance Domain
@@
TS data is commonly stored in spreadsheets and databases, which are prone to changes due to acts of omission (e.g., negligence, data-entry errors) or commission (e.g., adversarial attacks, sabotage). Omission errors are most common (Panko and Halverson, 1996). Tools like Microsoft Excel and Google Sheets are widely used for data collection and analysis, allowing end-user programming (Birch et al., 2018). However, over 90% of spreadsheets contain errors due to issues like incorrect formulae, leading to multi-billion dollar losses (Pak-Lok POON and TANG, 2024). Adversarial attacks are also increasing in data stores and AI models for tasks like forecasting. Malicious agents target ML models in financial institutions for financial gain, exploiting the limited robustness of deep-learning architectures commonly used in NLP. (Karim et al., 2019) explore both black-box and white-box attacks in time-series forecasting task. (Oregi et al., 2018) revealed the vulnerability of distance-based classifiers. (Rathore et al., 2020) examined various adversarial attacks on time series classifiers. TSFool (Li et al., 2022) introduced a multi-objective black-box attack to craft imperceptible adversarial time series to fool RNN classifiers. (Nehemya et al., 2021) highlights the vulnerability of algorithmic trading systems to real-time adversarial attacks using imperceptible perturbations, highlighting the need for mitigation strategies. (Fursov et al., 2021) examines adversarial attacks on deep-learning models in financial transaction records, revealing significant vulnerabilities. However, these works do not explore multi-modal models or multi-modal attacks. Our work addresses several perturbations applicable to uni-modal and multi-modal models, using both common data errors and attacks to measure the robustness of TSFM.
2.3. Robustness Testing of Time-series Forecasting Models
(Gallagher et al., 2022) examines the impact of different attacks on the performance of CNN model used for time series classification. (Pialla et al., 2023) introduces a stealthier attack using the Smooth Gradient Method (SGM) for time series and measures the effectiveness of the attack. While (Pialla et al., 2023) focuses on measuring the smoothness of the attacks, our work quantifies their impact on the models in addition to the biases they create in models’ predictions. (Govindarajulu et al., 2023) adapt attacks from the computer vision domain to create targeted adversarial attacks. They examine the impact of the proposed targeted attacks versus untargeted attacks using statistical measures. All these works measure the models’ performances under perturbations using statistical methods but do not measure the isolated impact of perturbations which is only possible through causal analysis. Furthermore, they do not consider any transformer-based or multi-modal models for evaluation.
2.4. Causal Analysis in Time-series Forecasting
(Moraffah et al., 2021) provides a review of the approaches used to compute treatment effects and also discusses causal discovery methods along with commonly used evaluation metrics and datasets. As the perturbations we introduce in this paper do not occur at the same timestep in each sample, the effect we are measuring can be considered as time-varying perturbation effect which is more complex to measure compared to time-invariant treatment effects. (Robins et al., 1999) measures the causal effect of one such time-varying exposure. However, they only consider binary outcomes. In our work, we deal with continuous outcomes and analyze the treatment effects of multiple treatments in the presence of confounders that are of interest in the time-forecasting domain.
2.5. Rating AI Models
@@
Several works have assessed and rated AI models for trustworthiness from a third-party perspective without access to training data. (Srivastava and Rossi, 2020) proposed a method to rate AI models for bias, specifically targeting gender bias in machine translators (Srivastava and Rossi, 2018), and used visualizations to communicate these ratings (Bernagozzi et al., 2021b). They conducted user studies on trust perception through visualizations (Bernagozzi et al., 2021a), but these lacked causal interpretation. (Lakkaraju et al., 2024) introduced a causal analysis approach to rate bias in sentiment analysis systems, extending it to assess their impact when used with translators (Lakkaraju et al., 2023). We extend this method to rate TSFM for robustness against perturbations. Causal analysis offers advantages over statistical analysis by determining accountability, aligning with humanistic values, and quantifying the direct influence of various attributes on forecasting accuracy.
3. Problem
3.1. Preliminaries
@@
Time Series Forecasting Let the time series be represented by {xt-n+1, xt-n+2, …., xt, xt+1, …, xt+d}, where each xt-n+i represents a value in time series, where $n$ is called the sliding window size and $d$ is the number of future values the model predicts. Let Xt == {xt-n+1, xt-n+2, …., xt}, and $\hat{Y_{t}}$ == $\hat{x}$t+1, $\hat{x}$t+2, …., $\hat{x}$t+d}, where $\hat{Y_{t}}=f(X_{t}$; $\theta$) for uni-modal TSFM, and in the case of multi-modal TSFM, $X_{t}$ includes a combination of numerical time-series values, time-series line plots, and time-frequency spectrograms. The function $f$ represents pre-trained TSFM with parameter $\theta$ that predicts the values for the next ‘d’ timesteps based on the values at previous $n$ timesteps. Let $Y_{t}$ denote the true values for the next ‘d’ timesteps. Let $S$ be the set of TSFM we want to rate. Let Rt be the residual for the sliding window [t + 1, t + d] and is computed by ($\hat{Y_{t}} -$ $Y_{t}$) at each timestep. Our rating method aims to highlight the worst-case scenario for the model. Therefore, we consider the maximum residual, denoted as $R^{max}_{t}$.
(a) Causal model $\mathcal{M}$ for TSFM.
@@
(b) Variants of the causal diagram in Figure 1(a) used to answer different research questions (RQs).
Figure 1. (a) Causal model $\mathcal{M}$ for TSFM. The validity of link ‘1’ depends on the data distribution ($P|Z$), while the validity of the links ‘2’ and ‘3’ are tested in our experiments. (b) Variants of $\mathcal{M}$ used to answer different research questions (RQs).
@@
Causal Model The causal model $\mathcal{M}$, is shown in Figure 1(a). Arrowheads indicate the causal direction from cause to effect. If Sensitive Attribute ($Z$) is a common cause for both Perturbation ($P$) and Residual ($R^{max}_{t}$), it introduces a spurious correlation between $P$ and $R^{max}_{t}$, known as the confounding effect, making $Z$ the confounder. The path from Perturbation to Residual through the confounder is called a backdoor path and is undesirable. Various backdoor adjustment techniques can remove the confounding effect (Xu and Gretton, 2022; Fang et al., 2024; Liu et al., 2021). The deconfounded distribution, after adjustment, is represented as $(R^{max}_{t}|do(P))$. The ‘do(.)’ operator in causal inference denotes an intervention to measure the causal effect of $P$ on the $R^{max}_{t}$. Solid red arrows with ‘?’ in Figure 1(a) denote the causal links tested in our experiments, while the dotted red arrow represents a potential causal link, depending on the distribution $(P|Z)$ across different values of $Z$.
3.2. Problem Formulation
@@
We aim to answer the following research questions (RQs) (with causal diagrams in Fig 1(b)) through our causal analysis when different perturbations denoted by $P$ = {0, 1, 2, 3} (or simply $P0,P1,$...) are applied to the input given to the set of TSFM $S$:
@@
RQ1: Does $Z$ affect $R^{max}_{t}$, even though $Z$ has no effect on $P$? That is, if perturbations are independent of the sensitive attribute, can the sensitive attribute still affect the model outcome, leading to statistical bias (i.e., lack of fairness)? (Perturbations are distributed uniformly with respect to $Z$; see Fig. 1(b), left).
@@
RQ2: Does $Z$ affect the relationship between $P$ and $R^{max}_{t}$ when $Z$ has an effect on $P$? That is, if the applied perturbations depend on the value of the sensitive attribute, would the sensitive attribute add a spurious (false) correlation between the perturbation and the outcome of a model leading to confounding bias? (Perturbations vary systematically with $Z$; see Fig. 1(b), middle).
@@
RQ3: Does $P$ affect $R^{max}_{t}$ when $Z$ may have an effect on $R^{max}_{t}$? That is, what is the impact of the perturbation on the outcome of a model when the sensitive attribute may still have an effect on the outcome of a model? (Same setup as RQ2, but focused on estimating the perturbation’s effect; see Fig. 1(b), right).
@@
RQ4: Does $P$ affect the accuracy of $S$? That is, do the perturbations affect the performance of the models’ accuracy? Causal analysis is not required to answer this question as we only need to compute appropriate accuracy metrics to assess how robust a model is against different perturbations.
4. Solution Approach